# Approximation rates for finite mixtures of location-scale models

### v2

@@ 1.1 Background and prior results
Finite mixture models appear ubiquitously throughout the statistics literature and their application and theory have been widely and comprehensively espoused in classic works such as Everitt & Hand, (1981), Titterington et al., (1985) McLachlan & Basford, (1988), Lindsay, (1995); McLachlan & Peel, (2000), and Frühwirth-Schnatter, (2006), along with more recent texts such as McNicholas, (2015), Ng et al., (2019), Robert et al., (2019), Chen, (2023), and Yao & Xiang, (2024), among many other sources. In application, finite mixtures are typically employed to perform one of two primary tasks: to provide a statistical model for a population that is hypothesized or known to have a heterogenous subpopulation structure, or to provide a flexible approximation to some unknown probability distribution. When the finite mixtures are employed for the latter task, the Gaussian mixture model is typically considered for its folkloric ability to provide accurate approximation typically encountered probability models (cf. Nguyen & McLachlan,, 2019 and references therein).
@@
When we restrict the approximand $f_{0}:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ to be a probability density function, and the approximator class ($\cal A$}, say) to be convex combinations of probability density kernels, the approximation problem requires different tools. When $f_{0}$ is within the the closure of $\cal A$}, and the class $\cal A$} is uniformly bounded in ${\cal L}_{p}$ norm, a Maurey–Barron–Jones type result of Donahue et al., (1997) provides an approximation theorem with known and optimal rate (see, e.g., Zeevi & Meir,, 1997 and Nguyen & McLachlan,, 2019 for example). However, when $f_{0}$ is not in the closure of $\cal A$} or when $\cal A$} is unbounded in ${\cal L}_{p}$ norm, the study of approximation demands more care.
@@ 1.2 Main contributions
In this work, we extend upon the qualitative results of Nguyen et al., (2020) and Nguyen et al., 2023b , by proving that it is also possible to obtain the ${\cal L}_{p}$ approximation rate in the number of mixture components (i.e., number of vertices of the approximating convex combination) when conducting approximation using location-scale mixture models when some smoothness assumption on $f_{0}$ hold. Particularly, for classes of targets $f_{0}$ in the Sobolev space ${\cal W}^{1,p}\left(\mathbb{R}^{d}\right)$ (Leoni,, 2017) or the fractional Sobolev spaces ${\cal W}^{s,p}\left(\mathbb{R}^{d}\right)$ ($s\in\left(0,1\right)$; Leoni,, 2023) with bounded norms, $f_{0}$ has smoothness coefficient $\alpha=1$ or $\alpha=s$, respectively. Then, for every $f_{0}$ whose Sobolev (or fractional Sobolev) norm is bounded, if $\cal A$} and $f_{0}$ satisfy appropriate integrability assumptions, then there exists sequences of location-scale mixtures $\left(f_{m}\right)_{m}\subset{\cal A}$ such that
@@
As an application of our result, we revisit the problem of adaptive least-squares density estimation considered in Klemelä, (2007) in a joint approximation and estimation setup, where the the expected squared ${\cal L}_{2}$ error between the least-squares estimator and the target $f_{0}$ is shown to converge at rate $n^{-1/2}$ when $f_{0}$ is in the closure of $\cal A$}, by taking estimator sequences with numbers of mixture components increasing at rate $n^{1/2}$ (see, also Klemela,, 2009, Ch. 19). We provide an analogous result for situations when $f_{0}$ is not in the closure of $\cal A$}, and demonstrate that that convergence rate of $n^{-\alpha/\left(2\alpha+d\right)}$ can be achieved by similarly using estimators whose components number increase at the same rate.
@@ 1.3 Related literature
Before we proceed to the body of our manuscript, we note that the literature on density estimation and approximation by finite mixture models is far broader and richer than what has been mentioned thus far. An incomplete list of such works in the area are as follows. For instance, the results of Nguyen et al., (2020) and Nguyen et al., 2023b has been extended to the mixture of experts setting of Jacobs et al., (1991) in Nguyen et al., (2021) and to the spherical manifold setting in Ng & Kwong, (2022). Schemes relating to weak approximations by finite mixture models appear in Dalal & Hall, (1983), Petrone & Veronese, (2010), and Lee & Lin, (2012), while best approximation and estimation results in various divergence measures are obtained by Ma et al., (2025) for the case where $f_{0}$ is in the closure of the Gaussian mixture models. Joint approximation and estimation results, analogous to Klemelä, (2007), have been investigated by Li & Barron, (1999), Rakhlin et al., (2005), and Dalalyan & Sebbar, (2018) in the Kullback–Leibler (KL) setting; by Genovese & Wasserman, (2000) in the Hellinger distance setting; and with a modified KL divergence in Chiu Chong et al., (2024). Beyond density estimation and approximation using adaptive least-squares estimators, there also exists a rich body of literature on parameter estimation via maximum likelihood estimation for mixture models (see e.g., Nguyen,, 2013, Shen et al.,, 2013, Ho & Nguyen, 2016a, , and Ho & Nguyen, 2016b, ) and for the related mixtures of experts class (see e.g., Jiang & Tanner,, 1999, Nguyen et al., 2023a, , Nguyen et al., 2024a, , Nguyen et al., 2024b, , and Nguyen et al., 2024c, ).
The rest of the manuscript proceeds as follows. In Section 2, we outline the notational conventions of the manuscript and provide some preliminary setup. In Section 3, we prove the main approximation theoretic result, and then in Section 4, we apply the approximation theorem to obtain the previewed least-squares estimation rate. A discussion of the work is provided in Section 5. The Appendix then contain the proof of the main result and collects some required technical results involved in our proofs.
@@ 4.1 Proof of Theorem 3
Let ${\cal F}\subset{\cal L}_{2}\cap{\cal P}$. If $f_{0}$ were known, then for each $\epsilon>0$, we can can approximate $f_{0}$ by some well-defined $f_{\epsilon}\in{\cal F}$, satisfying
@@
$\int_{\mathbb{R}^{d}}\hat{f}_{n,\epsilon}\left(x\right)\text{d}x-2P_{n}\hat{f}_{n,\epsilon}\leq\inf_{f\in{\cal F}}\left[\int_{\mathbb{R}^{d}}f^{2}\left(x\right)\text{d}x-2P_{n}f\right]+\epsilon\text{.}$ (3)
@@
$\int_{\mathbb{R}^{d}}\hat{f}_{m,n,\epsilon}\left(x\right)\text{d}x-2P_{n}\hat{f}_{m,n,\epsilon}$ $\leq\inf_{f\in\text{co}_{m}\left({\cal P}_{\varphi,\nu}\right)}\left[\int_{\mathbb{R}^{d}}f^{2}\left(x\right)\text{d}x-2P_{n}f\right]+\epsilon$
@@ 5.1 Vapnik–Chervonenkis classes
We note that the assumption that the VC dimension of the class ${\cal P}_{\varphi}$ be finite in Theorem 3 is not restrictive. In fact, a very large class of functions on $\mathbb{R}^{d}$ can be proved to have finite VC dimension via the fact that if the class of sets ${\cal C}=\left\{\mathbb{G}_{f}\text{ | }f\in\mathcal{F}\right\}$, where $\mathbb{G}_{f}=\left\{\left(x,t\right)\in\mathbb{R}^{d}\times\mathbb{R}\text{ | }t\leq f\left(x\right)\right\}$, is a uniformly definable with respect to an o-minimal (order-minimal) structure in the model theoretic sense, then $\textsc{vc}\left({\cal F}\right)<\infty$ (cf., Van den Dries,, 1998, Ch. 5 and Vidyasagar,, 2003, Thm. 10.11).
@@
$x\mapsto\varphi\left(x\right)=\left(2\pi\right)^{-d/2}\exp\left(\frac{1}{2}\left\|x\right\|_{2}^{2}\right)$

### v3

@@ 1.1 Background and prior results
Finite mixture models appear ubiquitously throughout the statistics literature and their application and theory have been widely and comprehensively espoused in classic works such as Everitt & Hand, (1981), Titterington et al., (1985) McLachlan & Basford, (1988), Lindsay, (1995); McLachlan & Peel, (2000), and Frühwirth-Schnatter, (2006), along with more recent texts such as McNicholas, (2015), Ng et al., (2019), Robert et al., (2019), Chen, (2023), and Yao & Xiang, (2024), among many other sources. In application, finite mixtures are typically employed to perform one of two primary tasks: to provide a statistical model for a population that is hypothesized or known to have a heterogenous subpopulation structure, or to provide a flexible approximation to some unknown probability distribution. When the finite mixtures are employed for the latter task, the Gaussian mixture model is typically considered for its folkloric ability to provide accurate approximations of typically encountered probability models (cf. Nguyen & McLachlan,, 2019 and references therein).
@@
When we restrict the approximand $f_{0}:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ to be a probability density function, and the approximator class ($\cal A$}, say) to be convex combinations of probability density kernels, the approximation problem requires different tools. When $f_{0}$ is within the closure of $\cal A$}, and the class $\cal A$} is uniformly bounded in ${\cal L}_{p}$ norm, a Maurey–Barron–Jones type result of Donahue et al., (1997) provides an approximation theorem with known and optimal rate (see, e.g., Zeevi & Meir,, 1997 and Nguyen & McLachlan,, 2019 for example). However, when $f_{0}$ is not in the closure of $\cal A$} or when $\cal A$} is unbounded in ${\cal L}_{p}$ norm, the study of approximation demands more care.
@@ 1.2 Main contributions
In this work, we extend upon the qualitative results of Nguyen et al., (2020) and Nguyen et al., 2023b , by proving that it is also possible to obtain the ${\cal L}_{p}$ approximation rate in the number of mixture components (i.e., number of vertices of the approximating convex combination) when conducting approximation using location-scale mixture models when some smoothness assumption on $f_{0}$ hold. Particularly, for classes of targets $f_{0}$ in the Sobolev space ${\cal W}^{1,p}\left(\mathbb{R}^{d}\right)$ (Leoni,, 2017) or the fractional Sobolev spaces ${\cal W}^{s,p}\left(\mathbb{R}^{d}\right)$ ($s\in\left(0,1\right)$; Leoni,, 2023) with bounded norms, $f_{0}$ has smoothness coefficient $\alpha=1$ or $\alpha=s$, respectively. Then, for every $f_{0}$ whose Sobolev (or fractional Sobolev) norm is bounded, if $\cal A$} and $f_{0}$ satisfy appropriate integrability assumptions, then there exist sequences of location-scale mixtures $\left(f_{m}\right)_{m}\subset{\cal A}$ such that
@@
As an application of our result, we revisit the problem of adaptive least-squares density estimation considered in Klemelä, (2007) in a joint approximation and estimation setup, where the the expected squared ${\cal L}_{2}$ error between the least-squares estimator and the target $f_{0}$ is shown to converge at rate $n^{-1/2}$ when $f_{0}$ is in the closure of $\cal A$}, by taking estimator sequences with numbers of mixture components increasing at rate $n^{1/2}$ (see, also Klemela,, 2009, Ch. 19). We provide an analogous result for situations when $f_{0}$ is not in the closure of $\cal A$}, and demonstrate that convergence rate of $n^{-\alpha/\left(2\alpha+d\right)}$ can be achieved by similarly using estimators whose components number increase at the same rate.
@@ 1.3 Related literature
Before we proceed to the body of our manuscript, we note that the literature on density estimation and approximation by finite mixture models is far broader and richer than what has been mentioned thus far. An incomplete list of such works in the area are as follows. For instance, the results of Nguyen et al., (2020) and Nguyen et al., 2023b have been extended to the mixture of experts setting of Jacobs et al., (1991) in Nguyen et al., (2021) and to the spherical manifold setting in Ng & Kwong, (2022). Schemes relating to weak approximations by finite mixture models appear in Dalal & Hall, (1983), Petrone & Veronese, (2010), and Lee & Lin, (2012), while best approximation and estimation results in various divergence measures are obtained by Ma et al., (2025) for the case where $f_{0}$ is in the closure of the Gaussian mixture models. Joint approximation and estimation results, analogous to Klemelä, (2007), have been investigated by Li & Barron, (1999), Rakhlin et al., (2005), and Dalalyan & Sebbar, (2018) in the Kullback–Leibler (KL) setting; by Genovese & Wasserman, (2000) in the Hellinger distance setting; and with a modified KL divergence in Chiu Chong et al., (2024). Beyond density estimation and approximation using adaptive least-squares estimators, there also exists a rich body of literature on parameter estimation via maximum likelihood estimation for mixture models (see e.g., Nguyen,, 2013, Shen et al.,, 2013, Ho & Nguyen, 2016a, , and Ho & Nguyen, 2016b, ) and for the related mixtures of experts class (see e.g., Jiang & Tanner,, 1999, Nguyen et al., 2023a, , Nguyen et al., 2024a, , Nguyen et al., 2024b, , and Nguyen et al., 2024c, ).
The rest of the manuscript proceeds as follows. In Section 2, we outline the notational conventions of the manuscript and provide some preliminary setup. In Section 3, we prove the main approximation theoretic result, and then in Section 4, we apply the approximation theorem to obtain the previewed least-squares estimation rate. A discussion of the work is provided in Section 5. The Appendix then contains the proof of the main result and collects some required technical results involved in our proofs.
@@ 4.1 Proof of Theorem 3
Let ${\cal F}\subset{\cal L}_{2}\cap{\cal P}$. If $f_{0}$ were known, then for each $\epsilon>0$, we can approximate $f_{0}$ by some well-defined $f_{\epsilon}\in{\cal F}$, satisfying
@@
$\int_{\mathbb{R}^{d}}\hat{f}_{n,\epsilon}^{2}\left(x\right)\text{d}x-2P_{n}\hat{f}_{n,\epsilon}\leq\inf_{f\in{\cal F}}\left[\int_{\mathbb{R}^{d}}f^{2}\left(x\right)\text{d}x-2P_{n}f\right]+\epsilon\text{.}$ (3)
@@
$\int_{\mathbb{R}^{d}}\hat{f}_{m,n,\epsilon}^{2}\left(x\right)\text{d}x-2P_{n}\hat{f}_{m,n,\epsilon}$ $\leq\inf_{f\in\text{co}_{m}\left({\cal P}_{\varphi,\nu}\right)}\left[\int_{\mathbb{R}^{d}}f^{2}\left(x\right)\text{d}x-2P_{n}f\right]+\epsilon$
@@ 5.1 Vapnik–Chervonenkis classes
We note that the assumption that the VC dimension of the class ${\cal P}_{\varphi}$ be finite in Theorem 3 is not restrictive. In fact, a very large class of functions on $\mathbb{R}^{d}$ can be proved to have finite VC dimension via the fact that if the class of sets ${\cal C}=\left\{\mathbb{G}_{f}\text{ | }f\in\mathcal{F}\right\}$, where $\mathbb{G}_{f}=\left\{\left(x,t\right)\in\mathbb{R}^{d}\times\mathbb{R}\text{ | }t\leq f\left(x\right)\right\}$, is uniformly definable with respect to an o-minimal (order-minimal) structure in the model theoretic sense, then $\textsc{vc}\left({\cal F}\right)<\infty$ (cf., Van den Dries,, 1998, Ch. 5 and Vidyasagar,, 2003, Thm. 10.11).
@@
$x\mapsto\varphi\left(x\right)=\left(2\pi\right)^{-d/2}\exp\left(-\frac{1}{2}\left\|x\right\|_{2}^{2}\right)$